\begin{thebibliography}{10}
  
\bibitem{Vafa}
M. Bershadsky, A. Johansen, V. Sadov and C. Vafa,
Topological reduction of 4D SYM to 2D $\s$-models,
Preprint, 1995.

\bibitem{Desale}
U. Desale and S. Ramanan,
Classification of vector bundles of rank 2 on hyperelliptic curves,
{\em Inventiones Math.}
{\bf 38} 1976, 161-185.

\bibitem{D1}
S. K. Donaldson,
Floer homology and algebraic geometry,
{\em Vector bundles in algebraic geometry,}
London Math. Soc. Lecture Notes Series, {\bf 208}
Cambridge University Press, Cambridge, 1995, 119-138.

\bibitem{DK}
S. K. Donaldson and P. B. Kronheimer,
{\em The geometry of 4-manifolds,} 
Oxford University Press, 1990.

\bibitem{Gottsche}
G. Ellingsrud and L. G\"ottsche,
Variation of moduli spaces and Donaldson invariants under change of
polarisation,
{\em Journal reine angew. Math.\/} 
{\bf 467} 1995, 1-49.

\bibitem{King}
A. D. King and P. E. Newstead,
On the cohomology ring of the moduli space of rank 2 vector bundles on
a curve,
Liverpool Preprint, 1994.

\bibitem{KM}
P. B. Kronheimer and T. S. Mrowka,
Embedded surfaces and the structure of Donaldson's polynomial invariants,
{\em Jour. Differential Geometry,}
{\bf 41} 1995, 573-734.

\bibitem{McDuff}
D. McDuff and D. A. Salamon,
J-holomorphic curves and quantum cohomology,
Preprint.

\bibitem{wall}
V. Mu\~noz,
Wall-crossing formulae for algebraic surfaces with $q >0$,
alg-geom/9709002.

\bibitem{Floer}
V. Mu\~noz,
Ring structure of the Floer cohomology of $\S \x \SS^1$,
dg-ga/9710029.

\bibitem{Piu1}
S. Piunikhin,
Quantum and Floer cohomology have the same ring structure,
MIT Preprint, 1994.

\bibitem{Piunikhin}
S. Piunikhin, D. Salamon and M. Schwarz,
Symplectic Floer-Donaldson theory and quantum cohomology,
Warwick Preprint, 1995.

\bibitem{Qin}
Z. Qin,
Moduli of stable sheaves on ruled surfaces and their Picard groups,
{\em Jour. Reine ange Math.\/}
{\bf 433} 1992, 201-219.

\bibitem{Ruan}
Y. Ruan,
Topological sigma model and Donaldson type invariants in Gromov theory,
{\em Duke Math. Jour.\/}
{\bf 83} 1996, 461-500.

\bibitem{RT}
Y. Ruan and G. Tian,
A mathematical theory of quantum cohomology,
{\em Jour. Diff. Geom.\/}
{\bf 42} 1995, 259-367.

\bibitem{Siebert}
B. Siebert,
An update on (small) quantum cohomology, 
Preprint, 1997.

\bibitem{ST}
B. Siebert and G. Tian,
Recursive relations for the cohomology ring of moduli spaces of stable
bundles,
{\em Proceedings of 3rd G\"okova Geometry-Topology Conference 1994.}

\bibitem{ST2}
B. Siebert and G. Tian,
On quantum cohomology rings of Fano manifolds and a formula of Vafa and
Intriligator,
submitted to Duke Math. Journal.

\bibitem{Thaddeus}
M. Thaddeus,
Conformal field theory and the cohomology of the moduli space of
stable bundles,
{\em Jour. Differential Geometry,}
{\bf 35} 1992, 131-150.

\end{thebibliography}

\documentstyle[12pt,amscd,amssymb]{amsart}

\newcommand{\ha}{\hat{\alpha}}
\newcommand{\hb}{\hat{\beta}}
\newcommand{\hg}{\hat{\gamma}}
\newcommand{\hR}{\hat{R}}
\newcommand{\hI}{\hat{I}}

\newcommand{\exseq}[3]{0 \ar #1 \ar #2 \ar #3 \ar 0}
\newcommand{\seq}[3]{{#1}_{#2}, \ldots, {#1}_{#3}}

\newcommand{\M}{M_{\Sigma}}
\newcommand{\Y}{\Sigma \times {\Bbb S}^1}
\newcommand{\X}{\Sigma \times {\Bbb P}^1}
\newcommand{\Spz}{\text{Sp}\, (2g,{\Bbb Z})}

\newcommand{\surj}{\twoheadrightarrow}
\newcommand{\inc}{\hookrightarrow}
\newcommand{\ar}{\rightarrow}

\newcommand{\x}{\times}
\newcommand{\ox}{\otimes}
\newcommand{\iso}{\cong}

\newcommand{\point}{\text{pt}}

\newcommand{\End}{\text{End}}
\newcommand{\Diff}{\text{Diff}}

\newcommand{\Gr}{\text{Gr}\,}

\newcommand{\coker}{\text{coker}}

\newcommand{\ch}{\text{ch}\:}
\newcommand{\Ext}{\text{Ext}}

\newcommand{\id}{\text{id}}

\newcommand{\cE}{{\cal E}}
\newcommand{\cF}{{\cal F}}

\newcommand{\cM}{{\cal M}}

\newcommand{\cN}{{\cal N}}
\newcommand{\cL}{{\cal L}}
\newcommand{\cO}{{\cal O}}

\newcommand{\cU}{{\cal U}}
\newcommand{\cV}{{\cal V}}

\renewcommand{\AA}{{\Bbb A}}
\newcommand{\CC}{{\Bbb C}}

\newcommand{\EE}{{\Bbb E}}

\newcommand{\PP}{{\Bbb P}}
\newcommand{\QQ}{{\Bbb Q}}

\newcommand{\SS}{{\Bbb S}}
\newcommand{\ZZ}{{\Bbb Z}}

\renewcommand{\a}{\alpha}
\renewcommand{\b}{\beta}

\newcommand{\g}{\gamma}

\newcommand{\f}{\epsilon}

\renewcommand{\l}{\lambda}

\newcommand{\s}{\sigma}

\renewcommand{\o}{\omega}
\newcommand{\p}{\phi}
\newcommand{\q}{\psi}

\newcommand{\z}{\zeta}

\renewcommand{\S}{\Sigma}

\renewcommand{\L}{\Lambda}
\renewcommand{\P}{\Phi}
\newcommand{\Q}{\Psi}

\newcommand{\frg}{{\frak g}}
\newcommand{\frM}{{\frak M}}

\theoremstyle{plain}
\begingroup
\theorembodyfont{\sl}
\newtheorem{thm}{Theorem}
\newtheorem{cor}[thm]{Corollary}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}

\endgroup

\theoremstyle{definition}
\newtheorem{defn}[thm]{Definition}

\theoremstyle{remark}
\newtheorem{rem}[thm]{Remark}
\newtheorem{ex}[thm]{Example}

\title[Quantum cohomology of moduli space of stable bundles]{First
  quantum correction for 
  the moduli space of stable bundles over a Riemann surface}

\author{Vicente Mu\~noz}
\address{Departamento de \'Algegra, Geometr\'{\i}a y Topolog\'{\i}a \\ Facultad
de Ciencias \\ Universidad de M\'alaga \\ 29071 M\'alaga \\ Spain}
\email{vmunoz@@agt.cie.uma.es}
\thanks{\hbox{$^*$}Supported by a grant from Ministerio de Educaci\'on y
Cultura \\ 1991 Mathematics Subject Classification. Primary: 57R57. Secondary:
58D27, 53C23, 57R57.}
\date{November, 1997}

\begin{document}

\maketitle

\begin{abstract}
  We compute some Gromov-Witten invariants of the moduli space $\M$ 
  of odd degree rank two
  stable vector bundles over a Riemann surface $\S$ of genus $g \geq 2$. 
  We thus find the first correction term for the quantum product of $\M$ and
  hence get the two leading terms of the relations satisfied by the 
  natural generators of the quantum cohomology of $\M$.
  Finally, we use this information to get a full description of the quantum
  cohomology of $\M$ when the genus of $\S$ is $g=3$. 
\end{abstract}

\section{Introduction}
\label{sec:intro}

Let $\S$ be a Riemann surface of genus $g \geq 2$ and let $\M$ denote the
moduli space
of flat $SO(3)$-connections with nontrivial second Stiefel-Whitney class $w_2$.

This is a smooth symplectic manifold
of dimension $6g-6$. Alternatively, we can consider $\S$ as a smooth complex
curve of
genus $g$. Fix a line bundle $\L$ on $\S$ of degree $1$, then 
$\M$ is the moduli space of  rank two 
stable vector bundles on $\S$ with determinant $\L$, which is a smooth
complex variety of complex dimension $3g-3$. The symplectic deformation class 
of $\M$ 
only depends on $g$ and not on the particular complex structure on $\S$.

The manifold $X=\M$ is a positive symplectic manifold with $\pi_2(X)=\ZZ$.
For such a manifold $X$, its quantum cohomology,
$QH^*(X)$, is well-defined 
(see~\cite{Ruan}~\cite{RT}~\cite{McDuff}~\cite{Piunikhin}).
As vector spaces, $QH^*(X)=H^*(X)$ (rational coefficients are understood),
but the multiplicative structure is different. 
Let $A$ denote the positive generator of $\pi_2(X)$, i.e. the
generator such that the symplectic form evaluated on $A$ is positive.
Let $N=c_1(X)[A] \in \ZZ_{>0}$. Then there is a natural $\ZZ/2N\ZZ$-grading 
for $QH^*(X)$, which 
comes from reducing the $\ZZ$-grading of $H^*(X)$. 
(For the case $X=\M$, $N=2$, so $QH^*(\M)$ is $\ZZ/4\ZZ$-graded).
The ring structure of 
$QH^*(X)$, called quantum multiplication, 
is a deformation of the usual cup product for $H^*(X)$. 
For $\a \in H^p(X)$, $\b \in H^q(X)$, we define the quantum 
product of $\a$ and $\b$ as 
$$
  \a \cdot \b =\sum_{d \geq 0} \P_{dA}(\a, \b),
$$
where $\P_{dA}(\a, \b) \in H^{p+q-2Nd} (X)$ is given by 
$<\P_{dA}(\a, \b),\g>=\Q^X_{dA}(\a, \b, \g)$, the Gromov-Witten invariant, for
all
$\g \in H^{\dim X-p-q+2Nd} (X)$. One has $\P_0(\a, \b)=\a \cup \b$. 
The other terms are
the correction terms and all live in lower degree parts of the
cohomology groups. It is a fact~\cite{RT} that the quantum product gives an
associative and graded commutative ring structure. 

To define the Gromov-Witten invariant, 
let $J$ be a generic almost complex
structure compatible with the symplectic form. 
Then for every $2$-homology class $dA$, $d \in \ZZ$, 
there is a moduli space $\cM_{dA}$ of pseudoholomorphic rational 
curves (with respect to $J$) 
$f : \PP^1 \ar X$ with $f_*[\PP^1] = dA$.
Note that $\cM_0=X$ and that $\cM_{dA}$ is empty for $d<0$.
For $d \geq 0$,
the dimension of $\cM_{dA}$ is $\dim X +2Nd$. 
This moduli space
$\cM_{dA}$ admits a natural compactification, $\overline{\cM}_{dA}$, called
the
Gromov-Uhlenbeck compactification~\cite{Ruan}~\cite[section 3]{RT}.
Consider now $r \geq 3$ different points $\seq{P}{1}{r} \in {\PP}^1$. Then we
have defined an evaluation map $ev: \cM_{dA} \ar X^r$
by $f \mapsto (f(P_1), \ldots, f(P_r))$. This map extends to 
$\overline{\cM}_{dA}$ and its image, $ev(\overline{\cM}_{dA})$, is a
pseudo-cycle~\cite{RT}. So for $\a_i \in
H^{p_i}(\M)$, $1 \leq i \leq r$, with $p_1+ \cdots +p_r = \dim X + 2Nd$, we 
choose generic cycles $A_i$, $1 \leq i \leq r$,
representatives of their Poincar\'e duals, and set
\begin{equation}
\label{eqn:qu1}
  \Q_{dA}^X(\seq{\a}{1}{r}) = <A_1\x\cdots\x A_r, [ev(\overline{\cM}_{dA})]>= 
  \# ev_{P_1}^*(A_1) \cap \cdots \cap ev_{P_r}^*(A_r),
\end{equation}
where $\#$ denotes count of points (with signs) and
$ev_{P_i}: \cM_{dA} \ar X$, $f \mapsto f(P_i)$.
This is a well-defined number and independent of the particular cycles.
Also, as the manifold $X$ is positive, 
$\coker L_f=H^1(\PP^1, f^* c_1(X))=0$, for all $f \in \cM_{dA}$.
By~\cite{Ruan}
the complex structure of $X$ is generic and we can use it to compute the
Gromov-Witten
invariants.

Also for $r \geq 2$, let $\a_i \in H^{p_i}(\M)$, $1 \leq i \leq r$, then
$$
  \a_1 \cdots \a_r =\sum_{d \geq 0} \P_{dA}(\a_1, \ldots ,\a_r),
$$ 
where the correction terms $\P_{dA}(\a_1, \ldots ,\a_r) \in
H^{p_1+\cdots +p_r - 2 N d}(X)$ are determined by 
$<\P_{dA}(\a_1, \ldots ,\a_r), \g> = \Q_{dA}^X(\a_1, \ldots ,\a_r, \g)$, for
any 
$\g \in H^{\dim X +2Nd-(p_1+\cdots +p_r)}(X)$.

Returning to our manifold $X=\M$, there is a classical conjecture relating
the quantum cohomology $QH^*(\M)$ and the instanton Floer cohomology of
$\Y$, $HF^*(\Y)$ (see~\cite{Floer}).
In~\cite{Vafa} a presentation of $QH^*(\M)$ was given using physical 
methods, and in~\cite{Floer} it was proved that such a presentation 
was a presentation of $HF^*(\Y)$ indeed. 
Siebert and Tian have a program~\cite{Siebert} to find the presentation
of $QH^*(\M)$, which goes through proving a recursion formula for the
Gromov-Witten invariants of $\M$ in terms of the genus $g$. This will
complete the proof of the conjectural isomorphism $QH^*(\M) \iso HF^*(\Y)$.

The purpose of this paper is two-fold. On the one hand we aim to compute
the Gromov-Witten invariants $\Q_A^{\M}$, relating them 
to the Donaldson invariants for the algebraic surface
$S=\S\x\PP^1$. This gives the first correction term of the quantum
product of $\M$ and hence the first two leading terms of the relations
satisfied by the generators of $QH^*(\M)$. In particular we obtain the
coefficient $c_g$ of~\cite[section 3.3]{Siebert}. 
On the other hand we infer a full presentation of $QH^*(\M)$ when the
genus of $\S$ is $g=3$ (recall that the case of genus $g=2$ was worked
out by Donaldson~\cite{D1}). 
This is the starting point of the induction
in~\cite{Siebert}.

The paper is organised as follows. In section~\ref{sec:ordinary} we review the

ordinary cohomology ring of $\M$. In section~\ref{sec:curves} the moduli space
of
lines in $\M$ is described, in order to compute the corresponding Gromov-Witten

invariants in section~\ref{sec:GW-inv}. In section~\ref{sec:quantum} we study
the
quantum cohomology of $\M$ and finally determine it completely in the case of
genus $g=3$ in section~\ref{sec:g3}.

\noindent {\em Acknowledgements:\/} I want to thank my D. Phil.\
supervisor Simon Donaldson, for his encouragement and invaluable help. 
Also I am very grateful to Bernd Siebert and Gang Tian for helpful
conversations and for letting me have a copy of their 
preprint~\cite{Siebert}.

\section{Classical cohomology ring of $\M$}
\label{sec:ordinary}

Let us recall the known description of the homology of
$\M$~\cite{King}~\cite{ST}~\cite{Floer}.
Let $\cU \ar \S \x \M$ be the universal bundle and consider the K\"unneth
decomposition of
\begin{equation}
   c_2(\End_0 \, \cU)=2 [\S] \ox \a + 4 \q -\b,
\label{eqn:qu2}
\end{equation}
with $\q=\sum \g_i \ox \q_i$, where 
$\{\seq{\g}{1}{2g}\}$ is a symplectic basis of 
$H^1(\S;\ZZ)$ with $\g_i \g_{i+g}=[\S]$ for $1 \leq i \leq g$ (also
$\{\g^{\#}_i\}$ will denote the dual basis for $H_1(\S;\ZZ)$). 
Here we can suppose without loss of generality that $c_1(\cU)=\L + \a$
(see~\cite{ST}).
In terms of the map $\mu: H_*(\S) \to H^{4-*}(\M)$,
given by $\mu(a)= -{1 \over 4} \,
p_1(\frg_{\cU}) /a $ (here $\frg_{\cU} \to \S \x \M$ is the associated
universal $SO(3)$-bundle, and $p_1(\frg_{\cU}) \in H^4(\S\x\M)$ its first
Pontrjagin class), we have
$$
   \left\{ \begin{array}{l} \a= 2\, \mu(\S) \in H^2
   \\ \q_i= \mu (\g_i^{\#}) \in H^3, \qquad 1\leq i \leq 2g
   \\ \b= - 4 \, \mu(x) \in H^4       
    \end{array} \right.
$$
where $x \in H_0(\S)$ is the class of the point, and $H^i=H^i(\M)$.
These elements generate $H^*(\M)$ as a 
ring~\cite{King}~\cite{Thaddeus}, and $\a$ is the positive generator of
$H^2(\M;\ZZ)$. We can rephrase this as saying that there exists an 
epimorphism 
\begin{equation}
  \AA(\S)= \QQ[\a,\b ]\ox \L(\seq{\q}{1}{2g}) \surj H^*(\M)
\label{eqn:qu3}
\end{equation}
(the notation $\AA(\S)$ follows that of Kronheimer and Mrowka~\cite{KM}).
The mapping class group $\Diff(\S)$ acts on $H^*(\M)$, with the action
factoring
through the action of $\Spz$ on $\{\q_i\}$. 
The invariant part, $H_I^*(\M)$, is generated by
$\a$, $\b$ and $\g=-2 \sum_{i=0}^g \q_i\q_{i+g}$. Then
there is an epimorphism
\begin{equation}
  \QQ[\a,\b,\g ]  \surj H^*_I(\M)
\label{eqn:qu4}
\end{equation}
which allows us to write
$$
   H_I^*(\M)= \QQ [\a, \b, \g]/I_g,
$$
where $I_g$ is the ideal of relations satisfied by $\a$, $\b$ and $\g$. 
Recall that $\deg(\a)=2$, $\deg(\b)=4$ and $\deg(\g)=6$. From~\cite{ST}, a
basis for $H_I^*(\M)$ is given by the monomials $\a^a\b^b\g^c$, with
$a+b+c<g$.
For
$0 \leq k \leq g$, the primitive component of $\L^k H^3$ is 
$$
   \L_0^k H^3 = \ker (\g^{g-k+1} : \L^k H^3 \ar \L^{2g-k+2} H^3).
$$
Then the $\Spz$-decomposition of $H^*(\M)$ is~\cite{King} 
$$
   H^*(\M)= \bigoplus_{k=0}^g \L_0^k H^3 \ox \QQ [\a, \b, \g]/I_{g-k}.
$$
\begin{prop}[\cite{ST}]
\label{prop:1}
  For $g=1$, let $q_1^1=\a$, $q_1^2=\b$, $q_1^3=\g$. Define recursively, for $g
\geq 1$,
$$
   \left\{ \begin{array}{l} q_{g+1}^1= \a q_g^1 +g^2 q_g^2
   \\ q_{g+1}^2 = \b q_g^1 + {2g \over g+1}  q_g^3
   \\ q_{g+1}^3 = \g  q_g^1
    \end{array} \right.
$$
  Then $I_g=(q_g^1, q_g^2, q_g^3)$, for all $g \geq 1$. Note that
  $\deg(q_g^1)=2g$, $\deg(q_g^2)=2g+2$ and $\deg(q_g^3)=2g+4$.
\end{prop}

\section{Holomorphic lines in $\M$}
\label{sec:curves}

In order to compute the Gromov-Witten invariants $\Q_A^{\M}$, we need
to describe the space of lines, i.e. rational 
curves in $\M$ representing the generator $A \in H_2(\M;\ZZ)$,
$$
 \cM_A=\{f: \PP^1 \ar \M / \text{$f$ holomorphic, } f_*[\PP^1]=A\}.
$$

Let us fix some notation. 
Let $J$ denote the Jacobian variety of $\S$ parametrising line bundles of
degree $0$ and
let $\cL \ar \S\x J$ be the universal line bundle. If $\{\g_i\}$ is
the basis of $H^1(\S)$ introduced in section~\ref{sec:ordinary}
then $c_1({\cL}) = \sum \g_i \ox \p_i \in H^1(\S) \ox H^1(J)$, where 
$\{\p_i\}$ is a symplectic basis for $H^1(J)$. Thus $c_1({\cL})^2 =
-2 [\S] \ox \o \in H^2(\S) \ox H^2(J)$, where $\o=\sum_{i=1}^g \p_i \wedge
\p_{i+g}$ is the natural symplectic form for $J$.

Consider now the algebraic surface $S=\X$. It has irregularity $q=g \geq 2$, 
geometric genus $p_g=0$ and canonical bundle $K \equiv -2\S+(2g-2){\PP}^1$.
Recall that $\L$ is a fixed line bundle of degree $1$ on $\S$.
Fix the line bundle $L=\L \ox \cO_{\PP^1}(1)$ on $S$ (we omit all
 pull-backs)
with $c_1= c_1(L) \equiv \PP^1+\S$, 
and put $c_2=1$. 
The ample cone of $S$ is $\{a\PP^1 +b \S \, / \, a,b >0 \}$.
Let $H_0$ be a polarisation close to $\PP^1$ in the ample cone
and $H$ be a polarisation close to $\S$, i.e.
$H= \S +t\PP^1$ with $t$ small.
We wish to study the moduli space $\frM= \frM_H(c_1,c_2)$ of $H$-stable
bundles
over $S$ with Chern classes $c_1$ and $c_2$. 

\begin{prop}
\label{prop:2}
  $\frM$ can be described as a bundle  $\PP^{2g-1} \ar \frM=
\PP(\cE_{\z}^{\vee})
  \ar J$, where $\cE_{\z}$ is a bundle on $J$ with $\ch \cE_{\z} = 2g + 8 \o$.
  So $\frM$ is compact, smooth and of the expected dimension $6g-2$.
  The universal bundle $\cV \ar S \x \frM$ is given by
$$
  0 \ar \cO_{\PP^1}(1) \ox \cL \ox \l \ar \cV \ar \L \ox \cL^{-1} \ar 0,
$$
  where $\l$ is the tautological line bundle for $\frM$.
\end{prop}

\begin{pf}
  For the polarisation $H_0$, the moduli space of $H_0$-stable 
  bundles with Chern classes
  $c_1, c_2$ is empty by~\cite{Qin}. Now for $p_1=-4c_2+c_1^2=-2$ 
  there is only one wall,
  determined by $\z \equiv -\PP^1 +\S$
  (here we fix $\z= 2\S- L=\S-c_1(\L)$ as a divisor), 
  so the moduli space of $H$-stable bundles with 
  Chern classes $c_1, c_2$ is obtained by crossing the wall as described
  in~\cite{wall}.
  First, note that the results in~\cite{wall} use the hypothesis of $-K$
  being 
  effective, but the arguments work equally well with the weaker assumption
  of $\z$ being a good wall~\cite[remark 1]{wall} (see also~\cite{Gottsche}
  for
  the case of $q=0$). In our case, $\z \equiv -\PP^1+\S$ is a good wall
  (i.e. 
  $\pm \z+K$ are both not effective) with $l_{\z}=0$. Now with the notations
  of~\cite{wall}, $F$ is a divisor such that $2F-L \equiv \z$, i.e. $F=\S$.
  Also
  $\cF \ar S \x J$ is the universal bundle parametrising divisors
  homologically
  equivalent to $F$, i.e. $\cF=\cL \ox \cO_{\PP^1}(1)$. Let
  $\pi: S \x J \ar J$ be
  the projection. Then $\frM=E_{\z}= \PP(\cE_{\z}^{\vee})$, where
$$
  \cE_{\z} = {\cE}\text{xt}^1_{\pi}(\cO(L-\cF),\cO(\cF))=  R^1\pi_* 
  (\cO(\z )\ox \cL^2).
$$
  Actually $\frM$ is exactly the set
  of bundles $E$ that can be written as extensions
$$
  \exseq{\cO_{\PP^1}(1) \ox L}{E}{\L \ox L^{-1}}
$$
  for a line bundle $L$ of degree $0$.
  The Chern character is~\cite[section 3]{wall} $\ch \cE_{\z} = 2g +
  e_{K-2\z}$,
  where $e_{\a}= -2 (\PP^1 \cdot \a) \o$ (the class $\S$ defined
  in~\cite[lemma 11]{wall}
  is $\PP^1$ in our case).
  Finally, the description of the universal bundle follows
  from~\cite[theorem 10]{wall}.
\end{pf}

\begin{prop}
\label{prop:3}
  There is a well defined map $\cM_A \ar \frM$.
\end{prop}

\begin{pf}
  Every line $f:{\PP}^1 \ar \M$ gives a bundle $E=(\id_{\S}\x f)^*\cU$ over
  $\X$
  by pulling-back the universal bundle ${\cU} \ar \S \x \M$. Then for any 
  $t \in \PP^1$, the bundle $E|_{\S \x t}$ is defined by $f(t)$.
  Now, by
  equation~\eqref{eqn:qu2}, $p_1(E)= p_1(\cU)[\S \x A]= -2\a [A] =-2$.
  Since $c_1(E) = (\id_{\S}\x f)^*c_1(\cU)=\L +\S$, it must be $c_2=1$.
  To see that $E$ is $H$-stable, consider any sub-line bundle $L \inc E$
  with $c_1(L) \equiv a\PP^1 +b\S$. Restricting to any $\S \x t \subset \X$ 
  and using the
  stability of $E|_{\S \x t}$, one gets $a \leq 0$. Then $c_1(L) \cdot \S < 
  {c_1(E)\cdot \S \over 2}$, which yields the $H$-stability of $E$
  (recall that $H$ is close to $\S$). So $E \in \frM$.
\end{pf}

Now define $N$ as the set of extensions on $\S$ of the form 
\begin{equation} 
  0 \ar L \ar E \ar \L \ox L^{-1} \ar 0,
\label{eqn:qu5}
\end{equation}
for $L$ a line bundle of degree $0$. Then the groups 
$\Ext^1( \L \ox L^{-1}, L) =H^1(L^2 \ox
\L^{-1})= H^0(L^{-2} \ox \L \ox K)$ are of constant dimension $g$. Moreover 
$H^0(L^{2} \ox \L^{-1})=0$, so the moduli space $N$ which
parametrises extensions like~\eqref{eqn:qu1} is given as 
$N=\PP(\cE^{\vee})$,
where $\cE= {\cE}{\hbox{xt}}^1_p(\L \ox {\cL}^{-1},
{\cL})=R^1 p_*({\cL}^{2} \ox
\L^{-1})$, $p:\S\x J \ar J$ the projection.
Then we have a fibration ${\PP}^{g-1} \ar N=\PP(\cE^{\vee}) \ar J$. 
The Chern character of $\cE$ is
\begin{equation} 
\begin{array}{rcl}
\ch({\cal E}) &=& \ch(R^1 p_*({\cL}^{2}\ox \L ^{-1})) =
             - \ch(p_{!}( {\cL}^2 \ox \L ^{-1}) ) = \\
             &=& -p_*( (\ch\,{\cL})^2 \> (\ch\,\L)^{-1} \> \text{Todd} \> 
              T_{\S}) = \\
             &=& -p_*((1+c_1({\cL})+{1 \over
               2}c_1({\cL})^2)^2(1-\L)(1-{1 \over 2}K)) = \\
             &=& -p_*(1-{1 \over 2}K+2c_1({\cL}) -4 \o
               \ox [\S]-\L) = g+4\o.
\end{array} \label{eqn:qu6}
\end{equation}
It is easy to check that 
all the bundles in $N$ are stable, so there is a well-defined map 
$$ 
 i: N \ar \M.
$$

Now we wish to construct the space of lines in $N$.
Note that $\pi_2(N)=\pi_2(\PP^{g-1})=\ZZ$, 
as there are no rational curves in $J$. Let
$L \in \pi_2(N)$ be the positive generator. We want to 
describe 
$$
   \cN_L=\{f: \PP^1 \ar N / \text{$f$ holomorphic, } f_*[\PP^1]=L\}.
$$
For the projective space $\PP^n$, the space of lines $H_1$ 
is the set of algebraic maps $f: {\PP}^1
\ar {\PP}^n$ of degree $1$. Such an $f$ has the form
$f[  x_0, x_1 ]=[x_0 u_0 +x_1 u_1]$, $[x_0,x_1] \in \PP^1$, 
where $u_0$, $u_1$ are linearly independent vectors 
in $\CC^{n+1}$. So
$$
  H_1= \PP(\{(u_0, u_1)/ \text{$u_0$, $u_1$ are linearly independent} \})
  \subset
  \PP((\CC^{\,n+1} \oplus \CC^{\,n+1})^{\vee}) =\PP^{2n+1}.
$$
The complement of $H_1$ is the image of $\PP^n \x \PP^1 \inc \PP^{2n+1}$,
$([u] , [x_0, x_1]) \mapsto [x_0 u, x_1 u]$, which is a smooth
$n$-codimensional
algebraic subvariety.
Now $\cN_L$ can be described as
\begin{equation}
  \begin{array}{ccccc}
     H_1     &\ar &  \cN_L   & \ar& J \\
  \bigcap    &    & \bigcap  &    & \| \\  
  \PP^{2g-1} &\ar & \PP((\cE \oplus \cE)^{\vee}) & \ar & J
  \end{array}
\label{eqn:qu7}
\end{equation}

\begin{rem}
\label{rem:4}
Note that $\cE_{\z}= R^1\pi_* (\cO(\z )\ox \cL^2)= R^1\pi_*(\cO_{\PP^1}(1) \ox
{\cL}^{2} \ox
\L^{-1})= H^0 (\cO_{\PP^1}(1)) \ox R^1p_*( {\cL}^{2} \ox \L^{-1})
= H^0 (\cO_{\PP^1}(1)) \ox \cE \iso \cE \oplus \cE$. So $\frM=\PP((\cE \oplus 
\cE)^{\vee})$, canonically.
\end{rem}

\begin{prop}
\label{prop:5}
The map $i:N \ar \M$ induces a map $i_*: \cN_L \ar \cM_A$. The composition
$\cN_L \ar \cM_A \ar \frM$ is the natural inclusion of~\eqref{eqn:qu7}.
\end{prop}

\begin{pf}
The first assertion is clear as $i$ is a holomorphic map.
For the second, consider the universal sheaf on $\S\x N$,
\begin{equation}
  0 \ar {\cL} \ox U \ar {\EE} \ar  \L \ox \cL^{-1} \ar 0,
\label{eqn:qu8}
\end{equation}
where $U=\cO_N(1)$ is the tautological bundle of
the fibre bundle $\PP^{g-1} \ar N \ar J$. 
Any element in $\cN_L$ is a line 
${\PP}^1 \inc N$, which must lie inside a single fibre $\PP^{g-1}$.
Restricting~\eqref{eqn:qu8} to this line, 
we have an extension 
$$
   0 \ar L \ox {\cO}_{{\PP}^1}(1) \ar E \ar \L \ox L^{-1}  \ar 0
$$
on $S=\X$, which is the image of the given element in $\frM$  
(here $L$ is the line bundle corresponding to the fibre 
in which ${\PP}^1$ sits). Now it is easy to check that the
map $\cN_L \ar \frM$ is the inclusion of~\eqref{eqn:qu7}.
\end{pf}

\begin{cor}
\label{cor:6}
$i_*$ is an isomorphism.
\end{cor}

\begin{pf}
By proposition~\ref{prop:5}, 
$i_*$ has to be an open immersion. The group $PGL(2,\CC)$ acts on both
spaces $\cN_L$ and $\cM_A$, and $i_*$ is equivariant.
The quotient $\cN_L/PGL(2,\CC)$ is compact, being
a fibration over the Jacobian with all the fibres the Grassmannian
$\Gr(\CC^{\,2},
\CC^{\,g-1})$, hence irreducible. As a consequence $i_*$ is an isomorphism.
\end{pf}

\begin{rem}
\label{rem:7}
Notice that the lines in $\M$ are all contained in the image of $N$,
which is of dimension $4g-2$ against $6g-6=\dim \M$.
They do not fill all of $\M$ as one would naively expect. 
\end{rem}

\section{Computation of $\Q_A^{\M}$}
\label{sec:GW-inv}

The manifold $N$ is positive with $\pi_2(N)=\ZZ$ and $L \in \pi_2(N)$ is
the positive generator. Under the map $i: N \ar \M$, 
we have $i_* L=A$. Now $\dim N= 4g-2$ and $c_1(N)[L]=c_1(\PP^{g-1})[L]=g$. So
quantum cohomology of $N$, $QH^*(N)$, is well-defined and
$\ZZ/2g\ZZ$-graded.
{}From corollary~\ref{cor:6}, it is straightforward to prove

\begin{lem}
\label{lem:8}
  For any $\a_i \in H^{p_i}(\M)$, $1 \leq i \leq r$, such that  $p_1 + \cdots +
p_r
  =6g-2$, it is
  $\Q_A^{\M}(\seq{\a}{1}{r})= \Q_L^N(\seq{i^*\a}{1}{r})$. $\quad \Box$
\end{lem}

It is therefore important to know the Gromov-Witten invariants of $N$, i.e.
its quantum cohomology.
{}From the universal bundle~\eqref{eqn:qu8}, we can read the first Pontrjagin
class $p_1( {\frg}_{\EE}) = -8  [\S]\ox\o + h^2- 2 [\S]\ox h
 + 4h\cdot c_1({\cL}) \in H^4(\S\x N)$, where $h=c_1(U)$ is the hyperplane
class. So on $N$ we have
\begin{equation}
  \left\{ \begin{array}{l}
  \a=2 \mu(\S) = 4\o + h \\
  \q_i=\mu(\g_i^{\#}) = - h \cdot \p_i \\
  \b= -4 \mu(x) =  h^2 
  \end{array} \right. 
\label{eqn:qu9}
\end{equation}
Let us remark that $h^2$ denotes ordinary cup product in $H^*(N)$,
a fact which will prove useful later. Now
let us compute the quantum cohomology ring of $N$. The cohomology of $J$ is
$H^*(J)= \L H_1$, where $H_1=H_1(\S)$. Now the fibre bundle 
description $\PP^{2g-1} \ar N=\PP(\cE^{\vee}) \ar J$
implies that the usual cohomology of $N$ is
$H^*(N)= \L H_1 [h]/<h^g+c_1h^{g-1}+ \cdots +c_g=0>$, where  $c_i=c_i(\cE)={4^i
 \over i!} \o^i$, from~\eqref{eqn:qu6}.
As the quantum cohomology has the same generators as the
usual cohomology and the relations are a deformation
of the usual relations~\cite{ST2}, it must be $h^g+c_1h^{g-1}+ \cdots +c_g=r$
in $QH^*(N)$, with
$r \in \QQ$. As in~\cite[example 8.5]{RT}, $r$ can be computed to be $1$. So
\begin{equation}
  QH^*(N)= \L H_1 [h]/<h^g+c_1h^{g-1}+ \cdots +c_g=1> .
\label{eqn:qu10}
\end{equation}

\begin{lem}
\label{lem:9}
  For any $s \in H^{2g-2i}(J)$, $0 \leq i \leq g$, denote by $s \in
  H^{2g-2i}(N)$ its pull-back to $N$ under the natural projection. 
  Then the quantum product $h^{2g-1+i} s$
  in $QH^*(N)$ has component in $H^{4g-2}(N)$ equal to ${(-8)^i\over i!}
  \o^i \wedge s$.
\end{lem}

\begin{pf}
  First note that for $s_1,s_2 \in H^*(J)$ such that their 
  cup product in $J$ is $s_1s_2=0$, then the quantum product $s_1s_2 \in
  QH^*(N)$ vanishes. This is so since every rational line in $N$ is 
  contained in a fibre of $\PP^{2g-1} \ar J \ar N$.

  Next note that $h^{g-1+i}s$ has component in $H^{4g-2}(N)$ equal to
  $s_i(\cE) \wedge s={(-4)^i\over i!} \o^i \wedge s$.
  Then multiply the standard relation~\eqref{eqn:qu9}
  by $h^{g-1+i}$ and work by induction on $i$. For $i=0$ we get
  $h^{2g-1}s=h^{g-1}s$ and the assertion is obvious. For $i>0$, 
$$
  h^{2g-1+i}s +h^{2g-2+i}c_1s+ \cdots +h^{2g-1}c_is=h^{g-1+i}s.
$$
  So the component of $h^{2g-1+i}s$ in $H^{4g-2}(N)$ is 
$$
  - \sum_{j=1}^i  {(-8)^{i-j}\over (i-j)!} \o^{i-j} c_j s +
  {(-4)^i\over i!} \o^i s =
  {(-8)^i\over i!} \o^i s
  - \sum_{j=0}^i  {(-8)^{i-j}\over (i-j)!}{4^j\over j!} \o^i s
  + {(-4)^i\over i!} \o^i s =   {(-8)^i\over i!} \o^i s.
$$
\end{pf}

\begin{lem}
\label{lem:10} 
Suppose $g>2$. Let $\a^a\b^b\q_{i_1}\cdots\q_{i_r} \in \AA(\S)$ 
have degree $6g-2$. Then 
$$
  \Q^N_L(\a, \stackrel{(a)}{\ldots}, \a, \b ,
  \stackrel{(b)}{\ldots} ,\b,\q_{i_1}, \ldots, \q_{i_r})=
  <(4\o + X)^a(X^2)^b \p_{i_1}\cdots\p_{i_r} X^r, [J]>,
$$
  evaluated on $J$, where $X^{2g-1+i}={(-8)^i  \over i!} \o^i \in H^*(J)$.
\end{lem}

\begin{pf}
  By definition the left hand side is the component in $H^{4g-2}(N)$
  of the quantum product $\a^a\b^b\q_{i_1}\cdots\q_{i_r} \in QH^*(N)$. 
  From~\eqref{eqn:qu9}, this quantum product is $(4\o + h)^a (h^2)^b
  (-h\p_{i_1})\cdots(-h\p_{i_r})$, 
  upon noting that when $g>2$, $\b=h^2$  
  as a quantum product as there are no quantum corrections 
  because of the degree.
  Note that $r$ is even, so the statement of the lemma follows from 
  lemma~\ref{lem:9}.
\end{pf}

Now we are in the position of relating the Gromov-Witten invariants 
$\Q_A^{\M}$ with the Donaldson invariants for $S=\S\x \PP^1$ 
(for definition of Donaldson invariants see~\cite{DK}~\cite{KM}).

\begin{thm}
\label{thm:11}
  Suppose $g>2$. Let $\a^a\b^b\q_{i_1}\cdots\q_{i_r} \in \AA(\S)$ have 
  degree $6g-2$. Then 
$$
  \Q^{\M}_A(\a, \stackrel{(a)}{\ldots}, \a, \b ,
  \stackrel{(b)}{\ldots} ,\b,\q_{i_1}, \ldots, \q_{i_r})=
  (-1)^{g-1} D^{c_1}_{S, H} ((2\S)^a(-4 \point)^b \g_{i_1}^{\#}
  \cdots\g_{i_r}^{\#}),
$$
  where $D^{c_1}_{S, H}$ stands for
  the Donaldson invariant of $S=\S\x\PP^1$ with 
  $w=c_1$ and polarisation $H$.
\end{thm}

\begin{pf}
  By definition, the right hand side is $\f_S(c_1)
  <\a^a\b^b\q_{i_1}\cdots\q_{i_r}, [\frM]>$, where 
  $\a =2\mu(\S) \in H^2(\frM)$, $\b =-4\mu(x) \in H^4(\frM)$,
  $\q_i =\mu(\g_i^{\#}) \in H^3(\frM)$.
  Here the factor $\f_S(c_1)=(-1)^{K_S c_1 +c_1^2
  \over 2}=(-1)^{g-1}$ compares the complex
  orientation of $\frM$ and its natural orientation as a moduli space of
  anti-self-dual connections~\cite{DK}. 
  By~\cite[theorem 10]{wall}, this is worked out to be 
  $(-1)^{g-1} <(4\o + X)^a(X^2)^b \p_{i_1}\cdots\p_{i_r} X^r,[J]>$, where 
  $X^{2g-1+i}=s_i(\cE_{\z})={(-8)^i  \over i!} \o^i$. Thus the 
  theorem follows from 
  lemmas~\ref{lem:8} and~\ref{lem:10}.
\end{pf}

\begin{rem}
\label{rem:12}
The formula in theorem~\ref{thm:11} is not right for $g=2$, as in
such case, the quantum product $h^2 \in QH^*(N)$ differs from $\b$ by
a quantum correction.
\end{rem}

\begin{rem}
\label{rem:13}
  Suppose $g \geq 2$ and let 
  $\a^a\b^b\q_{i_1}\cdots\q_{i_r} \in \AA(\S)$ have 
  degree $6g-6$. Then 
\begin{eqnarray*}
  \Q^{\M}_0(\a, \stackrel{(a)}{\ldots}, \a, \b ,
  \stackrel{(b)}{\ldots} ,\b,\q_{i_1}, \ldots, \q_{i_r}) &=&
  \f_S(\PP^1) < \a^a\b^b\q_{i_1}\cdots\q_{i_r},[\M]> = \\
  &=& - D^{\PP^1}_{S, H} ((2\S)^a(-4 \point)^b
  \g_{i_1}^{\#} \cdots\g_{i_r}^{\#}),
\end{eqnarray*}
as the moduli space of anti-self-dual connections on $S$ of dimension
$6g-6$ is $\M$.
\end{rem}

\section{Quantum cohomology ring of $\M$}
\label{sec:quantum}

The action of the mapping class group on $\M$ is symplectic, so
the quantum product restricts to the invariant part of the cohomology, thus
having defined $QH^*_I(\M)$ (see~\cite[section 3.1]{Siebert}). 
To give a description of it, let us define
  \begin{equation}
   \left\{ \begin{array}{l} \ha=\a \\ \hb=\b + r_g \\ 
   \hg =-2 \sum \q_i\q_{i+g} \end{array} \right. 
  \label{eqn:qu12} 
  \end{equation}
where $r_g \in \QQ$, $g \geq 1$ (to be determined shortly), 
and $\hg$ is given using the quantum product. 
It might happen that $\hg=\g +s_g \a$, $s_g \in \QQ$.
The need of introducing this quantum corrections
in the generators was noticed already in~\cite{Vafa}.

As a consequence of the description of $H_I^*(\M)$ given in
section~\ref{sec:ordinary}, the results in~\cite{ST2} imply that
there is a presentation
$$
  QH^*_I(\M)= \QQ [\ha, \hb, \hg]/J_g,
$$
where the ideal $J_g$ is generated by three elements $Q_g^1$, $Q_g^2$ and
$Q_g^3$
which are deformations graded mod $4$ of
$q_g^1$, $q_g^2$ and $q_g^3$, respectively (see~\cite{Siebert}).
This means that $J_g=(Q^1_g,Q^2_g,Q^3_g)$ for
\begin{equation}
\label{eqn:qu11}
   Q_g^i= \sum_{j \geq 0} Q_{g,j}^i,
\end{equation}
where $\deg(Q_{g,j}^i)=\deg(q_g^i)-4j$, $j \geq 0$, and $Q_{g,0}^i=q_g^i$.
There is still one source of possible ambiguity coming from adding a scalar
multiple
of $Q_g^1$ to $Q_g^3$. To avoid this, we require the coefficient of $\ha^g$ in
$Q_g^3$ to be zero.
Recall the main result from~\cite{Floer}.

\begin{thm}[\cite{Floer}]
\label{thm:14}
  Define $R^1_0=1$, $R^2_0=0$, $R^3_0=0$ and then recursively, for all 
  $g \geq 1$,
  $$
   \left\{ \begin{array}{l} R_{g+1}^1 = \a R_g^1 + g^2 R_g^2
   \\ R_{g+1}^2 = (\b+(-1)^{g+1}8) R_g^1 + {2g \over g+1}  R_g^3
   \\ R_{g+1}^3 =  \g  R_g^1
    \end{array} \right.
  $$
  Then the invariant part of the instanton Floer cohomology of $\Y$ is
  $HF^*(\Y)_I =\QQ[\a,\b,\g]/(R^1_g,R^2_g,R^3_g)$. $R^1_g$, $R^2_g$, $R^3_g$
are 
  uniquely
  determined by the conditions that the leading term of $R^i_g$ is $q^i_g$,
  $i=1,2,3$ and that
  the coefficient of $\a^g$ in $R^3_g$ is zero.
\end{thm}

Accounting for the difference in signs between theorem~\ref{thm:11} 
and remark~\ref{rem:13}, the
prospective generators of $J_g$ are defined as follows.

\begin{defn}
\label{def:15}
  For $i=1,2,3$ and $g \geq 1$, set
  $$
   \hR^i_g(\ha,\hb,\hg)= (\sqrt{-1})^{-\deg q^i_g}
   R^i_g(\sqrt{-1}^{\,g}\ha,\sqrt{-1}^{\,2g}\hb,
   \sqrt{-1}^{\,3g}\hg),
  $$
  i.e. when $g$ is even, $\hR^i_g=R^i_g$ and
  when $g$ is odd, $\hR^i_g$ is obtained from $R^i_g$ by changing the
  sign of the homogeneous components of degrees $\deg(q^i_g)-4-8j$, 
  $j\geq 0$.
\end{defn}

Thus we expect that $Q_g^i=\hR_g^i$, $i=1,2,3$ (i.e. $J_g=(\hat
 R_g^1,\hR_g^2,\hR_g^3)$), for $g \geq 1$
 (see~\cite{Vafa}~\cite{Siebert}). Let us review the known cases.

\begin{ex}
\label{ex:16}
  For $g=1$, we set $r_1=-8$, so that  
  $\ha=\a$, $\hb=\b - 8$, $\hg =\g$. Then the ideal of relations
  is generated by $\hR^1_1= \ha$, $\hR^2_1=\hb+ 8$
  and $\hR^3_1=\hat
  \g$. The correction $r_1=-8$ is arranged in such a way that things work. 
\end{ex}

\begin{ex}
\label{ex:17}
  For $g=2$, the quantum cohomology ring $QH^*(\M)$ has been computed by
  Donaldson~\cite{D1}, using an explicit description of $\M$ as the
  intersection of
  two quadrics in $\PP^5$. Let $h_2$, $h_4$ and $h_6$ be the integral 
  generators
  of $QH^2(\M)$, $QH^4(\M)$ and $QH^6(\M)$, respectively.
  Then, with our notations, $\a=h_2$, $\b = -4 h_4$ and
  $\g = 4 h_6$ (see~\cite{Vafa}). The computations in~\cite{D1} yield
  $\hg =\g -4 \a$. Now we set $r_2=4$, i.e. $\hb=\b + 4$. It is now
  easy to check that the relations found in~\cite{D1} can be 
  translated to 
  the relations $\hR_2^1= \ha^2+\hb-8$, $\hR_2^2= 
  (\hb+8) \ha +\hg$ and $\hR_2^3= \ha\hg$ 
  for $QH_I^*(\M)$.

  The artificially introduced term $r_2=4$ is due to the same phenomenon
  which
  causes the failure of lemma~\ref{lem:10} for $g=2$, i.e. the quantum 
  product $h^2$ differs from $\b$ in~\eqref{eqn:qu9}
  (defined with the cup product) because of a quantum correction in 
  $QH^*(N)$ which appears when $g=2$.
\end{ex}

In the general case we have

\begin{thm}
\label{thm:main}
  Let $g \geq 3$. Put $r_g=0$, so that $\ha=\a$ and $\hb=\b$. Then we can
write
$$
  \left\{
  \begin{array}{l} Q^1_g = \hR^1_g +f^1_g \\
  Q^2_g = \hR^2_g +f^2_g \\
  Q^3_g = \hR^3_g +f^3_g \end{array} \right.
$$
  where $\deg(f^i_g) \leq \deg(q^i_g) -8$, $i=1,2,3$.
\end{thm}

\begin{pf}
  As in equation~\eqref{eqn:qu11}, we can write $\hR_g^i= \sum_{j \geq 0}
\hR_{g,j}^i$,
  where $\deg(\hR_{g,j}^i)=\deg(q_g^i)-4j$, $j \geq 0$ (and analogously for
$R_g^i$).
  Clearly, $\hR_{g,0}^i=q_g^i=
  Q_{g,0}^i$, for $i=1,2,3$. We want to check that 
  $Q_{g,1}^i=\hR_{g,1}^i= (-1)^g R_{g,1}^i$. Pick any $z=
  \a^a\b^b\q_{i_1}\cdots\q_{i_r} \in \AA(\S)$ of degree $6g-2-\deg(q^i_g)$.
  By theorem~\ref{thm:14}, $\p^w(\S \x
  D^2,R^i_g)=0$ (see~\cite{Floer} for notations), so $D^{(w,\S)}_{S,H} (R^i_g
z)=0$, i.e.
  $$
   D^{\PP^1}_{S,H} (R^i_{g,1} z) + D^{c_1}_{S,H} (R^i_{g,0} z)=0.
  $$
  From theorem~\ref{thm:11} and remark~\ref{rem:13} this is translated as 
  the component in $H^{6g-6}(\M)$ of the quantum product
  $-R^i_{g,1} z + (-1)^{g-1}R^i_{g,0} z \in QH^*(\M)$ vanishing. 
  On the other hand,
  by definition $Q_g^i z =0 \in QH^*(\M)$, so the component in 
  $H^{6g-6}(\M)$ of the quantum product $Q^i_{g,1} z + Q^i_{g,0} 
  z \in QH^*(\M)$ is zero. Thus $<Q^i_{g,1}, z>= <(-1)^g R^i_{g,1}, z>$, for
any $z$
  of degree $6g-2-\deg(q^i_g)$, and hence $Q^i_{g,1} \equiv (-1)^g
  R^i_{g,1} \pmod{I_g}$ ($I_g$ is the ideal defined in
section~\ref{sec:ordinary}). 
  This gives the required equality $Q_{g,1}^i=\hR_{g,1}^i$
  (in the case $i=3$ we have to use the vanishing of the coefficient of
$\ha^g$
  for both $Q_g^3$ and $R_g^3$).
\end{pf}

\section{The case of genus $g=3$}
\label{sec:g3}

It is natural to ask to what extent the first quantum correction 
determine the
full structure of the quantum cohomology of $\M$. In~\cite{D1}, Donaldson
finds the first quantum correction for $\M$ when the genus of $\S$ is 
$g=2$ and proves that this determines the quantum product. In this section
we are going to check that this also happens for $g=3$, finding thus the
quantum cohomology of the moduli space of stable bundles over a Riemann
surface of genus $g=3$.

\begin{prop}
\label{prop:19}
  Let $\S$ have genus 
  $g=3$. Then $Q^1_3=\hR^1_3$, $Q^2_3=\hR^2_3$ and $Q^3_3=\hR^3_3$, i.e.
  $$
    QH^*_I(\M) = \QQ[\ha,\hb,\hg]/(\hR^1_3, \hR^2_3, \hR^3_3).
  $$
\end{prop}

\begin{pf}
  Theorem~\ref{thm:main} says that
  $$
  \left\{
  \begin{array}{l} Q^1_g = \hR^1_g = \ha(\ha^2 +\hb +8) + 4(\ha\hb -8\ha +\hg)
\\
  Q^2_g = \hR^2_g +x = (\hb+8)(\ha^2 +\hb +8) +{4 \over 3}\ha\hg +x \\
  Q^3_g = \hR^3_g +y \ha = \hg (\ha^2 +\hb +8)+y \ha
  \end{array} \right.
  $$
  where $x, y \in \QQ$. The main tool that we shall use is the nilpotency of
$\hg$. 
  Actually, from its definition~\eqref{eqn:qu12}, $\hg^4=0$ in $QH^*(\M)$.

  Suppose $y \neq 0$. Then the third relation implies $\ha^4=0$. Multiplying
the
  first relation by $\ha$ we get $5\ha^2\hb + 24\ha^2 +4 \ha\hg=0$. This is a
relation
  of degree $2g+2=8$. Thus it must be a multiple of $Q^2_g$, which is not.
This
  contradiction implies $y=0$.
 
  Let us see $x =0$. It is easy to check that
  $$
  \hg^3= {\hg^2 \over 4} \hR^1_3 -{3\hg(\hb-8) \over 4}\hR^2_3 + 
  {3(\hb+8)(\hb-8) -\ha\hg \over 4} \hR^3_3.
  $$
  The relations above imply that $\hg^3= -{3 \over 4} \hg(\hb-8) x$ in
$QH^*(\M)$.
  So $\hg^2(\hb-8)=0$. Also, the third relation (with $y=0$) gives $\hg\ha^2 =
  -\hg(\hb+8)$. Thus $\hg^2\ha^2=-16 \hg^2$ and $\hg^2\hb= 8\hg^2$.
  Multiply the first relation by $\hg^2$ to get $\hg^3=0$. As $\hg(\hb-8) \neq
0$
  (because this is not a multiple of the only relation of degree $2g=6$), it
must be
  $x=0$.
\end{pf}

\begin{cor}
\label{cor:20}
  Let $\S$ be a Riemann surface of genus $g=3$. Then 
  $$
   QH^*(\M)= \bigoplus_{k=0}^{g-1} \L_0^k H^3 \ox \QQ [\ha, \hb,
\hg]/\hI_{g-k}
  $$
  where we put $\hI_r=(
  R^1_r(\sqrt{-1}^{\,g}\ha,\sqrt{-1}^{\,2g}\hb,\sqrt{-1}^{\,3g}\hg), 
  R^2_r(\sqrt{-1}^{\,g}\ha,\sqrt{-1}^{\,2g}\hb,\sqrt{-1}^{\,3g}\hg),
\allowbreak
  R^3_r(\sqrt{-1}^{\,g}\ha,\sqrt{-1}^{\,2g}\hb,\sqrt{-1}^{\,3g}\hg))$, $1 \leq
r 
  \leq g$.
\end{cor}

\begin{pf}
  This is an easy consequence of the former proposition 
  and~\cite[lemma 7]{Floer}, noting that the hypothesis
  only has to be checked up to $g=3$. One has to be careful with the
  exponents of $\sqrt{-1}$ everywhere.
\end{pf}

\end{document}